OpenStudy (anonymous):

System of inequalities {y-x+2 Will someone go step by step with me and help me understand how to solve this math problem...plz.

OpenStudy (anonymous):

@snowblaze

OpenStudy (anonymous):

oh um i dont know though i will tag people

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

@lochana

OpenStudy (anonymous):

@Ddos_Dragon

OpenStudy (lochana):

yes. I can show you how it works.

OpenStudy (lochana):

first we draw graphs for each inequality. y = x -1 y = -x+2 

OpenStudy (lochana):

okay?

OpenStudy (lochana):

do you have any questions with that part?

OpenStudy (anonymous):

Now, how did you get the plot for x?

OpenStudy (lochana):

mm.. you need to find couple of points by putting random x values. then you can draw a line through them at the end.

OpenStudy (lochana):

I tell you how to draw graphs after this problem, but can you understand they represent y= x -1 and y = -x +2

OpenStudy (lochana):

but make sure that I didn't draw y < x -1 and y> -x +2 . I just graphed y = x -1 and y = -x +2

OpenStudy (anonymous):

Oh okay, so what step would take place after that?

OpenStudy (lochana):

yes. this is the valid points for y < x -1. notice darkened space.